MICROWAVE SPECTRA AND STRUCTURE OF N_{2} CO_{2}

\underline{DANIEL J. FROHMAN}, EDWIN S. CONTRERAS, ROSS S. FIRESTONE, STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459; WILLIAM KLEMPERER, Department of Chemistry and Chemical Biology, Harvard University, Cambridge, MA 02138.

We present here the FTMW spectra of the various isotopologues of the weakly-bound complex of carbon dioxide with the most abundant molecule in the atmosphere, nitrogen. The structure of the complex along with the inversion of the N_{2} (or lack of it) will be discussed. The molecule is T-shaped, with the OCO forming the cross of the T, a structure deduced from a previous rotationally-resolved infrared experiment.^{a} A significant wide-amplitude bending motion of the N_{2} is deduced from the values of the (nearly identical) nuclear quadrupole coupling constants of the nitrogen nuclei. The spectroscopic results will be compared with ab-initio calculations. We will examine the consequences of N_{2}-CO_{2} formation upon atmospheric opacity.

aM. A. Walsh, T. R. Dyke, B. J. Howard, J. Mol. Struct. 189, 111-120 (1988)